# Rigidity of proper quasi-homogeneous domains in positive flag manifolds

### v1

@@ Abstract.
We show that, inside the Shilov boundary of any given Hermitian symmetric space of tube type, there is, up to isomorphism, only one proper domain whose action by its automorphism group is cocompact. This gives a classification of all closed proper manifolds locally modelled on such Shilov boundaries, and provides a positive answer, in the case of flag manifolds admitting a $\Theta$-positive structure, to a rigidity question of Limbeek and Zimmer.
@@ 1.1. Divisible convex sets in real projective geometry
Convex projective geometry generalizes real hyperbolic geometry. The objects of study are the open sets $\Omega\subset\mathbb{P}(\mathbb{R}^{n})$ that are properly convex, meaning they are bounded and convex in an affine chart. The group of elements of $\operatorname{PGL}(n,\mathbb{R})$ preserving $\Omega$ is called the automorphism group of $\Omega$ and is denoted by $\mathsf{Aut}(\Omega)$. The set $\Omega$ is divisible if there exists a discrete subgroup of $\mathsf{Aut}(\Omega)$ that acts cocompactly on $\Omega$, and quasi-homogeneous if $\mathsf{Aut}(\Omega)$ acts cocompactly on $\Omega$. Note that divisibility implies quasi-homogeneity.
Properly convex open subsets of the projective space that are divisible are called divisible convex sets and admit a rich theory, which was initiated in the 1960’s with work of Benzecri [Ben60]. This theory has since then been developed by numerous authors (see e.g. [Vin65, Gol90, CLT15]), in particular by Benoist in the early 2000’s [Ben00, Ben03, Ben05, Ben06]. The strictly convex case is well understood, while the nonstrictly convex case is still in expansion [Isl, CLM20, Zim23, Bla]. See [Ben08] for a survey on this theory and more references.
@@
Even though it has recently been proven that divisible convex sets in the projective space are subject to a certain rigidity [Zim23], the diversity of nonsymmetric examples highlights the importance of general results concerning them, beyond those on cocompact actions on Riemannian symmetric spaces.
@@ 1.2. Flag manifolds and rigidity
One could also ask the same question, replacing “divisible” with “quasi-homogeneous”. A positive answer in the quasi-homogeneous case would imply a positive answer to Question 1.1.
First note that the study of proper quasi-homogeneous domains reduces to the case where $G$ is simple, by the following fact.
Fact 1.2 ([Zim18a, Thm 1.7]).
Let $G$ be a semisimple Lie group $G$ with trivial center and no compact factors, and write $G=G_{1}\times\cdots\times G_{k}$ where $G_{i}$ is a noncompact simple Lie group for all $1\leq i\leq k$. For any parabolic subgroup $P$ of $G$, there exist parabolic subgroups $P_{i}\leq G_{i}$ such that $P=P_{1}\times\cdots\times P_{k}$. Now let $\Omega\subset G/P$ be a proper quasi-homogeneous domain. Then there exist proper quasi-homogeneous domains $\Omega_{i}\subset G_{i}/P_{i}$ such that $\Omega=\Omega_{1}\times\cdots\times\Omega_{k}$.
@@
Let $G$ be a noncompact real simple Lie group and $P\leq G$ a nonmaximal proper parabolic subgroup. Then there are no proper quasi-homogeneous domains in $G/P$.
@@
In simultaneous joint work with A. Chalumeau [CG24], we give a positive answer to Question 1.1 when $G/P$ is the boundary of the pseudo-hyperbolic space $\mathbb{H}^{p,q}$ for $p\geq 2$ and $q\geq 1$ — also called the Einstein universe of signature $(p-1,q)$. The special case where $q=1$ admits a different proof than the general case, using causality in the Einstein universe of signature $(n-1,1)$. In the present paper, we generalize this causality approach to Shilov boundaries of Hermitian symmetric spaces of tube type (see the next section). These flag manifolds appear naturally in several contexts, such as Euclidean Jordan algebras and complex analysis (see e.g. [FK94]), or $\Theta$-positivity and higher Teichmüller Theory [GW16]. Here the proper parabolic subgroup $P$ is maximal, so Fact 1.3 does not apply. For any two transverse points $p,q\in G/P$, the open set consisting of all points of $G/P$ which are transverse to both $p$ and $q$ has several connected components, two of which are proper symmetric domains called diamonds (see Definition 4.6). The main result of the present paper (Theorem 1.4 below) is a positive answer to Question 1.1: we prove that in our context of Shilov boundaries there is, up to the action of $G$, only one proper quasi-homogeneous domain (hence only one proper divisible domain) in $G/P$, namely a diamond; in particular, it is symmetric.
@@
Let $G$ be a simple Lie group of Hermitian type of tube type and of real rank $r\geq 1$ (the complete list of the corresponding Lie algebras is given in Table 1). Let $\operatorname{\Delta}$ be set of simple restricted roots of $G$ and $\alpha_{r}\in\operatorname{\Delta}$ be the unique long root. Then the flag manifold $\mathscr{F}_{\{\alpha_{r}\}}=G/P_{\{\alpha_{r}\}}$ defined by $\{\alpha_{r}\}$ (see Section 2.2) is the Shilov boundary of the symmetric space $\mathbb{X}_{G}$ of $G$, denoted by $\mathbf{Sb}(G)$. Recall that the automorphism group $\mathsf{Aut}(\Omega)$ of a domain $\Omega\subset\mathbf{Sb}(G)$ is the group of all elements of $G$ that leave $\Omega$ invariant; we say that $\Omega$ is quasi-homogeneous if $\mathsf{Aut}(\Omega)$ acts cocompactly on $\Omega$.
@@
Let $G$ be a simple Lie group of Hermitian type of tube type. Then any proper quasi-homogeneous domain of $\mathbf{Sb}(G)$ is a diamond.
@@
More generally, if $G$ is a semisimple Lie group, Hermitian of tube type, with trivial center and no compact factors, then one can write $G=G_{1}\times\cdots\times G_{k}$, where $G_{i}$ is a noncompact simple Lie group, Hermitian of tube type, for all $1\leq i\leq k$. Then the Shilov boundary $\mathbf{Sb}(G)$ of $G$ is the flag manifold $\mathbf{Sb}(G_{1})\times\cdots\times\mathbf{Sb}(G_{k})$. By Fact 1.2 and Theorem 1.4, one directly gets:
@@
Let $G$ be a semisimple Lie group of Hermitian type of tube type with trivial center and no compact factors, and write $G=G_{1}\times\cdots\times G_{k}$, where $G_{i}$ is a noncompact simple Lie group of Hermitian type of tube type for all $1\leq i\leq k$. Then for any proper quasi-homogeneous domain $\Omega\subset\mathbf{Sb}(G)$, there are diamonds $D_{i}\subset\mathbf{Sb}(G_{i})$ for $1\leq i\leq k$, such that
@@ 1.4. $\Theta$-positive structures
Guichard–Wienhard generalized these two notions of total positivity and causality with their notion of $\Theta$-positive structure, where $\Theta$ is a subset of the simple roots of a semisimple Lie group $G$. They listed all pairs $(G,\Theta)$ such that $G$ admits a $\Theta$-positive structure [GW16, GW22]. The pairs $(G,\{\alpha_{r}\})$, where $G$ is a simple Lie group of Hermitian type of tube type and $\alpha_{r}$ is the simple root of $G$ defining the Shilov boundary $\mathbf{Sb}(G)$, constitute the only family of their list where the set $\Theta$ is a singleton (i.e. where the proper parabolic subgroup defined by $\Theta$ is maximal). Hence Theorem 1.4 and Fact 1.3 complete the classification of proper quasi-homogeneous domains in flag manifolds admitting a $\Theta$-positive structure:
@@
    If $|\Theta|=1$, then $G$ is Hermitian of tube type and $\mathscr{F}_{\Theta}=\mathbf{Sb}(G)$ admits exactly one proper quasi-homogeneous domain (up to the action of $G$), which is a diamond.
@@
    If $|\Theta|\geq 2$, then there does not exist any proper quasi-homogeneous domain in $\mathscr{F}_{\Theta}$.
Here we denote by $\mathscr{F}_{\Theta}$ the flag manifold defined by $(G,\Theta)$ (see Section 2.3). Again, by Fact 1.2, Question 1.1 has a positive answer for flag manifolds $G/P_{\Theta}$ with a $\Theta$-positive structure, where $G$ is a noncompact semisimple (not necessarily simple) Lie group and $\Theta$ a subset of the simple roots of $G$.
@@ 1.6. Comparison of two invariant metrics
In the classical theory of divisible convex sets, the Hilbert metric is a powerful tool to understand the automorphism group of a properly convex domain $\Omega$, due to its $\mathsf{Aut}(\Omega)$-invariance. For a proper quasi-homogeneous domain $\Omega$, the Hilbert metric allows one to relate the geometry of $\partial\Omega$ to the dynamics of $\mathsf{Aut}(\Omega)$. Namely, in this case any extremal point is “visual” (in the sense of Section 7.1):
@@
One idea in this paper is to adapt some arguments and results from convex projective geometry, in particular Fact 1.8, to the context of a flag manifold $\mathscr{F}_{\Theta}$ different from the real projective space. This requires constructing an $\mathsf{Aut}(\Omega)$-invariant metric on a proper domain $\Omega\subset\mathscr{F}_{\Theta}$ that gives enough information about the structure of $\partial\Omega$. Definition (2) above of the Hilbert metric already admits generalizations to arbitrary flag manifolds $\mathscr{F}_{\Theta}$ [Zim18a], called Caratheodory metrics and defined from irreducible $\Theta$-proximal representations of $G$; see Section 6.2.1. These metrics turn out to be proper (i.e. closed balls are compact) as soon as $\Omega$ is dually convex (see Section 4.1.2 and the proof of [Zim18a, Thm 9.1]). However, in our case, these metrics do not provide enough information about $\partial\Omega$ to prove Theorem 1.4. Instead, we use the Kobayashi metric, which generalizes Definition (1) above and is denoted by $K_{\Omega}$ in this paper. Contrary to those of Caratheodory metrics, the definition of the Kobayashi metric requires more structure on $\mathscr{F}_{\Theta}$: namely, the existence of photons satisfying some conditions of invariance (Lemmas 5.4 and 5.5) and of abundance (Lemma 6.2). These properties are satisfied for Shilov boundaries of irreducible Hermitian symmetric spaces of tube type, but not in arbitrary flag manifolds.
@@ 1.7.2. Diamonds
In Section 5, we define a photon-generating action as an action of $\operatorname{SL}_{2}(\mathbb{R})$ on $\mathbf{Sb}(G)$ which is conjugate to the one induced by the $\mathfrak{sl}_{2}$-triple associated with the last simple root of $\mathfrak{g}$. A photon is then the unique closed orbit in $\mathbf{Sb}(G)$ of a photon-generating action. It is a topological circle whose properties are similar to those of projective lines in $\mathbb{P}(\mathbb{R}^{n})$ (Lemmas 5.4 and 5.5) and allow us to define and analyze the Kobayashi metric (see Section 6), and to understand the boundary of a proper domain in $\mathbf{Sb}(G)$ (see Section 7).
@@ 1.7.4. The Kobayashi metric
A substantial part of Section 6 (more precisely, Sections 6.2.2 and 6.2.3) is devoted to proving Proposition 1.9. The complex roots of an intersection polynomial describe the intersection of complexifications of a photon and of a generic maximal proper Schubert subvariety of $\mathbf{Sb}(G)$; since this intersection is a singleton (Lemma 6.9), the intersection polynomials are split (Corollary 6.10). This observation allows us to compare two cross ratios, namely the one appearing in the definition of the Kobayashi metric and the other in that of the Caratheodory metric; see Lemma 6.11, from which follows Proposition 1.9.
@@
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper quasi-homogeneous domain. In Section 7, two families of points of $\partial\Omega$ are introduced: the visual points, defined by a dynamical property, and the $\mathscr{R}$-extremal points (see Definition 7.3), defined by a geometric one. The definition of $\mathscr{R}$-extremal points and that of the Kobayashi metric both involve photons, which allows us to prove an analogue of Fact 1.8, following the strategy of [LZ19]: namely, $\mathscr{R}$-extremal points are visual (Proposition 7.4).
By quasi-homogeneity, the domain $\Omega$ is dually convex [Zim18a, Thm 9.1], that is, for any $p\in\partial\Omega$ there is a supporting hypersurface to $\Omega$ at $p$ of the form $\operatorname{Z}_{z}$ (see (1.1)); see Section 4.1.2. Lemma 7.1 expresses a stronger geometric property for a visual point $p\in\partial\Omega$: this supporting hypersurface $\operatorname{Z}_{z}$ can be taken to be $\operatorname{Z}_{p}$ itself. This lemma, applied to two strongly $\mathscr{R}$-extremal points $p_{0},q_{0}\in\partial\Omega$ (see Definition 7.3 and Lemma 7.10), implies that $\Omega$ is contained in the diamond $\mathbf{D}(p_{0},q_{0})$ (Section 8).
@@
In Section 8, with the notation of Section 1.7.5 above, we prove that $\Omega=\mathbf{D}(p_{0},q_{0})$. The key point is the inclusion $\mathsf{Aut}(\Omega)\leq\mathsf{Aut}(\mathbf{D}(p_{0},q_{0}))$, which implies, by quasi-homogeneity, that $\Omega$ is closed in (and hence equal to) the diamond $\mathbf{D}(p_{0},q_{0})$. This inclusion holds because any automorphism of $\Omega$ preserves the pair $\{p_{0},q_{0}\}$, which is proved in Proposition 8.2, and essentially characterizes $\mathbf{D}(p_{0},q_{0})$ (Fact 4.8).

### v2

# Rigidity of proper almost-homogeneous domains in positive flag manifolds

@@ Abstract.
We show that, inside the Shilov boundary of any given Hermitian symmetric space of tube type, there is, up to isomorphism, only one proper domain such that every point on its boundary belongs to the closure of an orbit under its automorphism group. This gives a classification of all closed proper manifolds locally modelled on such Shilov boundaries, and provides a positive answer, in the case of flag manifolds admitting a $\Theta$-positive structure, to a rigidity question of W. van Limbeek and A. Zimmer.
@@ 1.1. Divisible convex sets in real projective geometry
Convex projective geometry generalizes real hyperbolic geometry. The objects of study are the open sets $\Omega\subset\mathbb{P}(\mathbb{R}^{n})$ that are properly convex, meaning they are bounded and convex in an affine chart. The group of elements of $\operatorname{PGL}(n,\mathbb{R})$ preserving $\Omega$ is called the automorphism group of $\Omega$ and is denoted by $\mathsf{Aut}(\Omega)$. The set $\Omega$ is divisible if there exists a discrete subgroup of $\mathsf{Aut}(\Omega)$ that acts cocompactly on $\Omega$, and quasi-homogeneous if $\mathsf{Aut}(\Omega)$ acts cocompactly on $\Omega$. It is said to be almost-homogeneous if its full orbital limit set, that is, the union of the accumulation points of all orbits of $\mathsf{Aut}(\Omega)$, is equal to its boundary. Note that divisibility implies quasi-homogeneity, which implies almost-homogeneity.
Properly convex open subsets of the projective space that are divisible are called divisible convex sets and admit a rich theory, which was initiated in the 1960’s with work of Benzecri [Ben60]. This theory has since then been developed by numerous authors (see e.g. [Vin65, Gol90, CLT15]), in particular by Benoist in the early 2000’s [Ben00, Ben03, Ben05, Ben06]. The strictly convex case is well understood, while the nonstrictly convex case is still in expansion [Isl25, CLM20, Zim23, Bla24]. See [Ben08] for a survey on this theory and more references.
@@
Even though it has recently been proved that divisible convex sets in the projective space are subject to a certain rigidity [Zim23], the diversity of nonsymmetric examples highlights the importance of general results concerning them, beyond those on cocompact actions on Riemannian symmetric spaces.
Other examples of properly convex sets arise by imposing weaker assumptions on the action of the automorphism group: certain properly convex sets in $\mathbb{P}(\mathbb{R}^{n})$ admit actions by discrete groups $\Gamma\leq\operatorname{PGL}(n,\mathbb{R})$ with finite covolume, without being divisible or even quasi-homogeneous. Such examples are almost-homogeneous [BM20].
@@ 1.2. Flag manifolds and rigidity
One could also ask the same question, replacing “divisible” with “almost-homogeneous”. A positive answer in the almost-homogeneous case would imply a positive answer to Question 1.1.
First note that the study of proper almost-homogeneous domains reduces to the case where $G$ is simple, by the following fact.
Lemma 1.2.
[Zim18a, Thm 1.7] Let $G$ be a semisimple Lie group $G$ with trivial center and no compact factors, and write $G=G_{1}\times\cdots\times G_{k}$ where $G_{i}$ is a noncompact simple Lie group for all $1\leq i\leq k$. For any parabolic subgroup $P$ of $G$, there exist parabolic subgroups $P_{i}\leq G_{i}$ such that $P=P_{1}\times\cdots\times P_{k}$. Now let $\Omega\subset G/P$ be a proper almost-homogeneous domain. Then there exist proper almost-homogeneous domains $\Omega_{i}\subset G_{i}/P_{i}$ such that $\Omega=\Omega_{1}\times\cdots\times\Omega_{k}$.
Although Lemma 1.2 is proved in [Zim18a, Thm 1.7] for quasi-homogeneous domains, their proof holds in the quasi-homogeneous case. We give it in Section 4.1.1.
@@
Let $G$ be a noncompact real simple Lie group and $P\leq G$ a nonmaximal proper parabolic subgroup. Then there are no proper almost-homogeneous domains in $G/P$.
As for Lemma 1.2, Fact 1.3 is proved in [Zim18a, Thm 1.5] for proper quasi-homogeneous domains, but their proof still holds for almost-homogeneous domains.
@@
In simultaneous joint work with A. Chalumeau [CG24], we give a positive answer to Question 1.1 when $G/P$ is the boundary of the pseudo-hyperbolic space $\mathbb{H}^{p,q}$ for $p\geq 2$ and $q\geq 1$ — also called the Einstein universe of signature $(p-1,q)$. The special case where $q=1$ admits a different proof than the general case, using causality in the Einstein universe of signature $(n-1,1)$. In Lemma 7.7 of the present paper, we generalize this causality approach to Shilov boundaries of Hermitian symmetric spaces of tube type (see the next section), in order to prove our main Theorem 1.4 below. These flag manifolds appear naturally in several contexts, such as Euclidean Jordan algebras and complex analysis (see e.g. [FK94]), or $\Theta$-positivity and higher Teichmüller Theory [GW16]. Here the proper parabolic subgroup $P$ is maximal, so Fact 1.3 does not apply. For any two transverse points $p,q\in G/P$, the open set consisting of all points of $G/P$ which are transverse to both $p$ and $q$ has several connected components, two of which are proper symmetric domains called diamonds (see Definition 4.6). The main result of the present paper (Theorem 1.4 below) is a positive answer to Question 1.1: we prove that in our context of Shilov boundaries there is, up to the action of $G$, only one proper almost-homogeneous domain (hence only one proper divisible domain) in $G/P$, namely a diamond; in particular, it is symmetric.
@@
Let $G$ be a simple Lie group of Hermitian type of tube type and of real rank $r\geq 1$ (the complete list of the corresponding Lie algebras is given in Table 1). Let $\operatorname{\Delta}$ be set of simple restricted roots of $G$ and $\alpha_{r}\in\operatorname{\Delta}$ be the unique long root. Then the flag manifold $\mathscr{F}_{\{\alpha_{r}\}}=G/P_{\{\alpha_{r}\}}$ defined by $\{\alpha_{r}\}$ (see Section 2.2) is the Shilov boundary of the symmetric space $\mathbb{X}_{G}$ of $G$, denoted by $\mathbf{Sb}(G)$. Recall that the automorphism group $\mathsf{Aut}(\Omega)$ of a domain $\Omega\subset\mathbf{Sb}(G)$ is the group of all elements of $G$ that leave $\Omega$ invariant; we say that $\Omega$ is almost-homogeneous if every element of $\partial\Omega$ lies in the closure of the orbit of an element of $\Omega$ under $\mathsf{Aut}(\Omega)$.
@@
Let $G$ be a simple Lie group of Hermitian type of tube type. Then every proper almost-homogeneous domain of $\mathbf{Sb}(G)$ is a diamond.
Since every divisible proper domain of $\mathbf{Sb}(G)$ is almost homogeneous, Theorem 1.4 gives a positive answer to Question 1.1. It also implies that, reciprocally, every almost-homogeneous domain of $\mathbf{Sb}(G)$ is divisible. As already mentionned in Section 1.1, this property does not hold in the projective space, as there exist properly convex domains of the projective space that are almost-homogeneous and not divisible, and not even quasi-homogeneous, see [BDL18].
@@
More generally, if $G$ is a semisimple Lie group, Hermitian of tube type, with trivial center and no compact factors, then one can write $G=G_{1}\times\cdots\times G_{k}$, where $G_{i}$ is a noncompact simple Lie group, Hermitian of tube type, for all $1\leq i\leq k$. Then the Shilov boundary $\mathbf{Sb}(G)$ of $G$ is the flag manifold $\mathbf{Sb}(G_{1})\times\cdots\times\mathbf{Sb}(G_{k})$. By Lemma 1.2 and Theorem 1.4, one directly gets:
@@
Let $G$ be a semisimple Lie group of Hermitian type of tube type with trivial center and no compact factors, and write $G=G_{1}\times\cdots\times G_{k}$, where $G_{i}$ is a noncompact simple Lie group of Hermitian type of tube type for all $1\leq i\leq k$. Then for any proper almost-homogeneous domain $\Omega\subset\mathbf{Sb}(G)$, there are diamonds $D_{i}\subset\mathbf{Sb}(G_{i})$ for $1\leq i\leq k$, such that
@@ 1.4. $\Theta$-positive structures
Guichard–Wienhard generalized these two notions of total positivity and causality with their notion of $\Theta$-positive structure, where $\Theta$ is a subset of the simple roots of a semisimple Lie group $G$. They listed all pairs $(G,\Theta)$ such that $G$ admits a $\Theta$-positive structure [GW16, GW22]. The pairs $(G,\{\alpha_{r}\})$, where $G$ is a simple Lie group of Hermitian type of tube type and $\alpha_{r}$ is the simple root of $G$ defining the Shilov boundary $\mathbf{Sb}(G)$, constitute the only family of their list where the set $\Theta$ is a singleton (i.e. where the proper parabolic subgroup defined by $\Theta$ is maximal). Hence Theorem 1.4 and Fact 1.3 complete the classification of proper almost-homogeneous domains in flag manifolds admitting a $\Theta$-positive structure:
@@
    If $|\Theta|=1$, then $G$ is Hermitian of tube type and $\mathscr{F}_{\Theta}=\mathbf{Sb}(G)$ admits exactly one proper almost-homogeneous domain (up to the action of $G$), which is a diamond.
@@
    If $|\Theta|\geq 2$, then there does not exist any proper almost-homogeneous domain in $\mathscr{F}_{\Theta}$.
Here we denote by $\mathscr{F}_{\Theta}$ the flag manifold defined by $(G,\Theta)$ (see Section 2.3). Again, by Lemma 1.2, Question 1.1 has a positive answer for flag manifolds $G/P_{\Theta}$ with a $\Theta$-positive structure, where $G$ is a noncompact semisimple (not necessarily simple) Lie group and $\Theta$ a subset of the simple roots of $G$.
@@ 1.6. Comparison of two invariant metrics
In the classical theory of divisible convex sets, the Hilbert metric is a powerful tool to understand the automorphism group of a properly convex domain $\Omega$, due to its $\mathsf{Aut}(\Omega)$-invariance. For a proper almost-homogeneous domain $\Omega$, the Hilbert metric allows one to relate the geometry of $\partial\Omega$ to the dynamics of $\mathsf{Aut}(\Omega)$:
@@
One idea in this paper is to adapt some arguments and results from convex projective geometry, in particular Fact 1.8, to the context of a flag manifold $\mathscr{F}_{\Theta}$ different from the real projective space. This requires constructing an $\mathsf{Aut}(\Omega)$-invariant metric on a proper domain $\Omega\subset\mathscr{F}_{\Theta}$ that gives enough information about the structure of $\partial\Omega$. Definition (2) above of the Hilbert metric already admits generalizations to arbitrary flag manifolds $\mathscr{F}_{\Theta}$ [Zim18a], called Caratheodory metrics and defined from irreducible $\Theta$-proximal representations of $G$; see Section 6.2.1. These metrics turn out to be proper (i.e. closed balls are compact) as soon as $\Omega$ is dually convex (see Section 4.1.2 and the proof of [Zim18a, Thm 9.1]). However, in our case, these metrics do not provide enough information about $\partial\Omega$ to prove Theorem 1.4. Instead, we use the Kobayashi metric, which generalizes Definition (1) above and is denoted by $K_{\Omega}$ in this paper. Contrary to those of Caratheodory metrics, the definition of the Kobayashi metric requires more structure on $\mathscr{F}_{\Theta}$: namely, the existence of photons satisfying some conditions of invariance (Lemmas 5.3 and 5.4) and of abundance (Lemma 6.2). These properties are satisfied for Shilov boundaries of irreducible Hermitian symmetric spaces of tube type, but not in arbitrary flag manifolds.
@@ 1.7.2. Diamonds
In Section 5, we define a photon-generating action as an action of $\operatorname{SL}_{2}(\mathbb{R})$ on $\mathbf{Sb}(G)$ which is conjugate to the one induced by the $\mathfrak{sl}_{2}$-triple associated with the last simple root of $\mathfrak{g}$. A photon is then the unique closed orbit in $\mathbf{Sb}(G)$ of a photon-generating action. It is a topological circle whose properties are similar to those of projective lines in $\mathbb{P}(\mathbb{R}^{n})$ (Lemmas 5.3 and 5.4) and allow us to define and analyze the Kobayashi metric (see Section 6), and to understand the boundary of a proper domain in $\mathbf{Sb}(G)$ (see Section 7).
@@ 1.7.4. The Kobayashi metric
A substantial part of Section 6 (more precisely, Sections 6.2.2 and 6.3) is devoted to proving Proposition 1.9. The complex roots of an intersection polynomial describe the intersection of complexifications of a photon and of a generic maximal proper Schubert subvariety of $\mathbf{Sb}(G)$; since this intersection is a singleton (Lemma 6.11), the intersection polynomials are split (Corollary 6.12). This observation allows us to compare two cross ratios, namely the one appearing in the definition of the Kobayashi metric and the other in that of the Caratheodory metric; see Lemma 6.13, from which follows Proposition 1.9.
@@
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper domain. In Section 7, we define the $\mathscr{R}$-extremal points (see Definition 7.1) of $\Omega$, which are analoguous to extremal points of a properly convex subset of the projective space. The definition of $\mathscr{R}$-extremal points and that of the Kobayashi metric both involve photons. Following the strategy of [LZ19], this allows us to prove an analogue of Fact 1.8; see Lemma 7.6, whose main consequence is a strong geometric property for $\mathscr{R}$-extremal points of an almost-homogeneous domain (see Theorem 7.2).
By almost-homogeneity, the domain $\Omega$ is dually convex (see [Zim18a, Thm 9.1] and Proposition 6.9), that is, for any $p\in\partial\Omega$ there is a supporting hypersurface to $\Omega$ at $p$ of the form $\operatorname{Z}_{z}$ (see (1.1)); see Section 4.1.2. Lemma 7.2 expresses a stronger geometric property for an extremal point $p\in\partial\Omega$: this supporting hypersurface $\operatorname{Z}_{z}$ can be taken to be $\operatorname{Z}_{p}$ itself. This lemma, applied to two strongly $\mathscr{R}$-extremal points $p_{0},q_{0}\in\partial\Omega$ (see Definition 7.1 and Lemma 7.7), implies that $\Omega$ is contained in the diamond $\mathbf{D}(p_{0},q_{0})$ (Section 8).
@@
In Section 8, with the notation of Section 1.7.5 above, we prove that $\Omega=\mathbf{D}(p_{0},q_{0})$. The key point is the inclusion $\mathsf{Aut}(\Omega)\leq\mathsf{Aut}(\mathbf{D}(p_{0},q_{0}))$, which implies, by almost-homogeneity, that $\Omega$ is closed in (and hence equal to) the diamond $\mathbf{D}(p_{0},q_{0})$. This inclusion holds because any automorphism of $\Omega$ preserves the pair $\{p_{0},q_{0}\}$, which is proved in Proposition 8.2, and essentially characterizes $\mathbf{D}(p_{0},q_{0})$ (Fact 4.8).